Find the tangent lines as directed to the parabola (x+2y)^2+2x-y-3=0 parallel to the line 4x+3y=2


1
Expert's answer
2021-11-30T14:35:03-0500

Let the tangent be x-4y=k => x= k+4y. Use this in the equation of the curve


"(4y+k)^2 +2(k+4y) -y -3 = 16y^2 +(8k+7)y +(k^2+2k-3)=0"


This should have equal roots so 

"(8k+7)^2\u20134(16)(k^2+2k-3)=9 or (112\u2013228)k+(49+192)=0"


k= 241/116